# Multi-language Unit Test Generation using LLMs

### v1

@@ Abstract.
Implementing automated unit tests is an important but time consuming activity in software development. Developers dedicate substantial time to writing tests for validating an application and preventing regressions. To support developers in this task, software engineering research over the past few decades has developed many techniques for automating unit test generation. However, despite this effort, usable tools exist for very few programming languages—mainly Java, C, and C# and, more recently, for Python. Moreover, studies have found that automatically generated tests suffer poor readability and often do not resemble developer-written tests. In this work, we present a rigorous investigation of how large language models (LLMs) can help bridge the gap. We describe a generic pipeline that incorporates static analysis to guide LLMs in generating compilable and high-coverage test cases. We illustrate how the pipeline can be applied to different programming languages, specifically Java and Python, and to complex software requiring environment mocking. We conducted a through empirical study to assess the quality of the generated tests in terms of coverage, mutation score, and test naturalness—evaluating them on standard as well as enterprise Java applications and a large Python benchmark. Our results demonstrate that LLM-based test generation, when guided by static analysis, can be competitive with, and even outperform, state-of-the-art test-generation techniques in coverage achieved while also producing considerably more natural test cases that developers find easy to read and understand. We also present the results of a user study, conducted with 161 professional developers, that highlights the naturalness characteristics of the tests generated by our approach.
unit test, java, python, llm
ccs: General and reference Empirical studiesccs: Computing methodologies Neural networks
1. Introduction
@@
Unit testing is a key activity in software development that serves as the first line of defense against the introduction of software bugs. Manually writing high-coverage and effective unit tests can be tedious and time consuming. To address this, many automated test generation techniques have been developed, aimed at reducing the cost of manual test suite development: over the last few decades, this research has a produced a variety of approaches based on symbolic analysis (e.g., (Visser et al., 2004; Cadar et al., 2008; Păsăreanu and Rungta, 2010; Godefroid et al., 2005; Sen et al., 2005; Xie et al., 2005; Tillmann and De Halleux, 2008)), search-based techniques (e.g., (Tonella, 2004; Harman and McMinn, 2010; McMinn, 2004; Fraser and Arcuri, 2011; Lin et al., 2021; Lukasczyk and Fraser, 2022)), random and adaptive-random techniques (e.g., (Pacheco et al., 2007; Ciupa et al., 2008; Chen et al., 2010; Lin et al., 2009; Arcuri and Briand, 2011; Lukasczyk et al., 2023)), and learning-based approaches (Durelli et al., 2019).
@@
These techniques have achieved considerable success in generating high-coverage test suites with good fault-detection ability, but they still have several key limitations—with respect to test readability, test scenarios covered, and test assertions created. Previous studies (e.g., (Fraser et al., 2015)) have shown that developers find automatically generated tests lacking in these characteristics, suffering poor readability and comprehensibility, covering uninteresting sequences, and containing trivial or ineffective assertions. Automatically generated tests are also known to contain anti-patterns or test smells (Panichella et al., 2020) and generally not perceived as being “natural” in the sense that they do not resemble the kinds of tests that developers write.
Recently, researchers have started to explore the potential of leveraging large language models (LLMs) to overcome these limitations of conventional test-generation techniques (Wang et al., 2024a). Given the inherent ability of LLMs in synthesizing natural-looking code (having been trained on massive amounts of human-written code), this is a promising research direction toward development of techniques that generate natural test cases. Early work in this area has focused on empirical studies of different LLMs for unit test generation using simple prompts or experimenting with various prompting strategies by including different types of context information suitable for test generation (e.g., (Tufano et al., 2020; Schäfer et al., 2023; Siddiq et al., 2024; Hashtroudi et al., 2023; Tang et al., 2024)), and investigation of feedback mechanisms for generating compilable code or achieving specific testing goals, such as increasing coverage or killing surviving mutants (e.g., (Schäfer et al., 2023; Ryan et al., 2024; Dakhel et al., 2023; Pizzorno and Berger, 2024)).
Although these explorations have demonstrated the potential of LLMs in generating natural-looking test cases, they have also highlighted drawbacks of LLM-based test generation, such as the LLM’s tendency to generate tests that often do not compile, achieve low coverage, contain testing anti-patterns, or are redundant in a generated test suite. Moreover, much of this early work explores LLMs off-the-shelf as tools for test generation without considering complete solutions that generate tests with some guarantees (e.g., tests that compile) as conventional tools, such as EvoSuite (evosuite, 2024), do. Finally, the empirical studies conducted so far are limited in the LLMs considered, the investigation of test naturalness, and the complexity of code targeted for test generation.
@@
Figure 1. Illustration of naturalness (in terms of test names, variable names, and assertions) and mocking in test cases generated by the LLM-assisted technique of aster (right) compared with tests generated by EvoSuite (Fraser and Arcuri, 2011) and CodaMosa (Lemieux et al., 2023) (left).
@@
In this work, we present a technique for performing LLM-assisted test generation guided by program analysis. The technique consists of preprocessing and postprocessing phases that wrap LLM interactions. The preprocessing phase performs static program analysis to compute relevant information to be included in LLM prompts for a given method under test. This ensures that the LLM prompt has sufficient context (similar to the information that a developer would use while writing test cases for a method) and increases the chances that it generates compilable and meaningful test cases. The postprocessing phase checks the generated tests for compilation and runtime errors and constructs new prompts aimed at fixing the errors. At the end of the test-repair process, the technique produces a set of passing test cases and a set of failing tests for the focal method (i.e., method under test). To increase code coverage, the technique includes a coverage-augmentation phase, in which prompts are crafted for instructing the LLM to generate test cases aimed at exercising uncovered lines of code.
@@
We implemented the technique in a tool, called aster, for two programming languages (PL), Java and Python, thus demonstrating the feasibility of building multi-lingual unit test generators with LLMs guided by lightweight program analysis. aster also incorporates mocking capability for Java unit test generation, which makes it applicable to applications that perform database operations, implement services, or use complex libraries. We present a generic approach for generation of test with mocks that is extensible to different library APIs. The approach performs static analysis to identify types in the focal class (i.e., the class under test) that need mocking and constructs a structured prompt, which includes a test skeleton with mocking setup, fixtures, and partial test code created and leaves the test-completion task for the LLM. With such prompts, LLMs are able, in many cases, to generate usable tests for scenarios involving Java Servlet API (javaservletspec, 2024), etc.
@@
We performed a comprehensive empirical evaluation of aster, using six general-purpose and code LLMs, to study the generated tests in terms of code coverage achieved, mutation score, and naturalness of test code; we also compared aster against state-of-the-art Java and Python unit test generators. Our results show that aster performs very competitive (+2.0%, -0.5%, and +5.1% line, branch, and method coverage) for Java SE applications while performing significantly better (+26.4%, +10.6%, and +18.5%, line, branch, and method coverage) for Java EE applications. For Python, aster has a high coverage gain compared to CodaMosa (+9.8%, +26.5%, and +22.5% line, branch, and method coverage). In terms of naturalness, both our automated evaluation and developers’ survey identified that aster-generated test cases are more natural compared to EvoSuite- and CodaMosa-generated tests and, surprisingly, better than even developer-written test cases. The results of our experiments and the naturalness evaluator are available in our artifact-source (asterartifact, 2024)
@@
  • Generic pipeline supporting multiple PLs. We present a generic LLM-assisted test-generation pipeline fueled by static analysis that can be extended for multiple PLs. We evaluate aster for Java and Python and compare it against state-of-the-art test generators (EvoSuite and CodaMosa).
  • Complex enterprise application. We evaluated aster on complex enterprise applications that often require special test generation capabilities such as mocking.
  • Multiple models. Our evaluation includes a wide variety of models ranging from closed-source (GPT) to open-source (Llama, Granite) as well as small and large models (8B to $>$1T).
@@
  • Evaluation of test naturalness. We performed an in-depth quantitative and qualitative analysis of test naturalness by developing an automated approach and conducting a survey of 161 professional developers.
2. Motivation
@@
The primary motivation for LLM-assisted test generation is to overcome the limitation of conventional test generation w.r.t lack of naturalness in the tests they create. By leveraging the LLM’s inherent ability of creating natural-looking code, we can generate more readable, comprehensible, and meaningful test cases. Another motivation for LLM-assisted test generation is to build multi-lingual unit test generators—leveraging the LLM’s understanding of the syntax and semantics of multiple PLs on which the models are typically trained. Building such test generators using conventional approaches (e.g., symbolic or evolutionary techniques) can be challenging and no such test generators exist. In contrast, with lightweight static analysis guiding LLM interactions, a multi-lingual unit test generator can be easily implemented. Finally, an LLM-assisted approach can also address test generation for scenarios requiring mocking.
@@
To illustrate these benefits, fig. 1 presents sample aster-generated test cases and tests generated by two conventional techniques: EvoSuite (Fraser and Arcuri, 2011) for Java and CodaMosa (Lemieux et al., 2023) for Python. The aster-generated test cases have more meaningful test names  1, variable names  2, and assertions  3 than the EvoSuite- or CodaMosa-generated tests. For instance, consider the test cases for Apache Commons CLI (commonscli, 2024). The variable storing the return value from flatten() is called flattenedArguments in the aster test—clearly capturing the meaning of the stored data—whereas the corresponding variable in the EvoSuite test is called stringArray1 (line 5), which captures simply the data structure type instead of any meaning of the stored data. Similarly, the Python test case generated by aster, shown in fig. 1(c), has meaningful test name and variable names. Moreover, the assertion in the test case is generated taking into account the expected transformation of the input by the focal method, converting the input hour value to minutes. fig. 1(b) shows a unit test case with mocking of library APIs generated for Apache Commons JXPath (commonsjxpath, 2024). The test mocks the behavior of the org.w3c.dom.Node library class  4. As we describe later, by creating a test skeleton containing mocking-related fixture via static analysis, our approach can generate usable test cases for complex applications requiring mocking of API calls.
3. Our Approach
@@
Figure 2 illustrates our LLM-based test-generation technique, guided by program analysis. The process comprises two phases, both utilizing stage-specific prompting: (1) Preprocessing, in which static analysis extracts context and template (fig. 3) constructs initial prompts for LLM-driven unit test generation in Java and Python and (2) Postprocessing, where the generated tests undergo code fixing for correctness and code coverage. The prompt template is modified with relevant content. This iterative approach leverages LLM capabilities while ensuring test suite robustness through systematic analysis and refinement.
3.1. Preprocessing
@@
The preprocessing phase performs static analysis of the application to gather all the context that an LLM might require to generate unit tests for Java and Python. The key objective of this stage is to gather all the necessary information pertaining to the focal method and its broader context within the application. This information is used to populate fields 1 and 2 of the prompt template shown in Figure 3. In this section, we discuss the preprocessing steps for Java and how those steps help construct a prompt for the LLM.
@@
Figure 3. Templates for composing prompts for test generation, test repair, and coverage augmentation.
@@
    Testing scope. The first step in the preprocessing phase identifies the testing scope given a Java class under test (i.e., the focal class) $f_{c}$. The testing scope lists the set of focal methods to be targeted for test generation. This set consists of (1) public, protected, and package-visibility methods declared in $f_{c}$ and (2) any inherited method implementations from an abstract super class of $f_{c}$. The set excludes inherited methods from a non-abstract super class as those methods are targeted for test generation in the context of their declaring class as the focal class. If $f_{c}$ is an abstract class, the testing scope consists of static methods, if any, declared in $f_{c}$. The focal class and method are used to populate part d of the test-generation prompt.
@@
    Relevant constructors. aster next identifies the relevant constructors for a focal method so that their signatures can be specified to the LLM, enabling it to create required objects for invoking the focal method. The relevant constructors for a focal method $m$ include the constructors of the focal class (if $m$ is a virtual method), along with the constructors of the each formal parameter type of $m$, considering application types only (i.e., ignoring library types). The analysis is done transitively for the formal parameter types of each identified constructor, thus ensuring that the LLM prompt includes comprehensive context information about how to create instances of application types that may be needed for testing $m$. The discovered constructors are used to populate part a of the test-generation prompt in fig. 3.
@@
    Relevant auxiliary methods. Accessor methods, which consist of getters and setters, provide a mechanism for reading and modifying internal object state, while preserving encapsulation. Our approach identifies setter methods in the focal class and in each formal parameter type of the focal method (if the type is an application class). Signatures of these setters are added to the LLM prompt to help with setting up suitable object states for invoking the focal method in the generated test case. For reading object state, our approach computes getters of the focal class and the return type of the focal method (limiting to application return types), and includes their signatures in the LLM prompt. This can help the LLM generate suitable assertion statements for verifying relevant object state after the focal method call—the assertions can check state of the receiver object of, or the object returned by, the focal method call. This information is used to fill section b of the test-generation prompt in fig. 3.
@@
    Private methods.  Private methods are inaccessible outside the class and therefore, to test them, we need to invoke them through a non-private method of the class. To facilitate this, we compute the class call graph of the focal class, identify call chains from non-private methods to private methods, and provide these call chains to the LLM to enable it to generate test cases that invoke private methods through externally visible methods. This information is used to populate part e of the test-generation prompt in fig. 3.
    Input : focal method $f_{m}$, focal class $f_{c}$, mockable APIs $\mathcal{T}_{apis}$
    Output : Candidate mockable Types $\mathcal{T}_{m}$, and mockable Fields $\mathcal{F}_{m}$
    1 Function IdentifyMockedFieldsAndParameterTypes:
           $\triangleright$ Identify mockable fields.
    2       $\mathcal{F}_{m},\mathcal{T}_{m}\leftarrow\emptyset,\emptyset$;
    3       foreach $f\in f_{c}.\text{fields}$ do
    4             if $f.\text{type}\in\mathcal{T}_{apis}$ then
    5                   $\mathcal{F}_{m}\leftarrow\mathcal{F}_{m}\cup\{f\}$;
    6
          $\triangleright$ Identify all mockable Types
    7       $T\leftarrow\{f_{c}\}\cup{\textnormal{{formalParamTypes}}(f_{m})}$;
    8       while $T\neq\emptyset$ do
    9             $t\leftarrow T.\text{pop()}$();
    10             $\mathcal{T}_{m}\leftarrow\mathcal{T}_{m}\cup\{t:t\in\mathcal{T}_{apis}\}$;
    11             foreach $c\in t.\text{constructors}$ do
    12                   $T\leftarrow T\cup\textnormal{{formalParamTypes}}(c)$;
    13
    14return $\mathcal{T}_{m}$, $\mathcal{F}_{m}$;
    15
    Input : focal method $f_{m}$, focal class $f_{c}$, and mockable types $\mathcal{T}_{apis}$.
    Output : Mockable constructor calls $\mathcal{M}_{c}$, static calls $\mathcal{M}_{s}$, API calls $\mathcal{M}_{a}$
    16 Function IdentifyMockingScope:
    17       $S,\mathcal{M}_{c},\mathcal{M}_{s},\mathcal{M}_{a}\leftarrow\emptyset,\emptyset,\emptyset,\emptyset$;
          $\triangleright$ Add focal method and constructor to scope
    18       $S\leftarrow S\cup\{f_{m},f_{c}.\textit{constructors}\}$;
          $\triangleright$ Add methods reachable from $f_{m}$ within $f_{c}$
    19       $S\leftarrow S\cup\{m\in f_{c}.\textit{methods}:\textnormal{{isReachable}}(f_{m},m)\}$;
          $\triangleright$ Add service entry class methods to scope
[…]
    21             $S\leftarrow S\cup\{m\in f_{c}.\textit{methods}:\textnormal{{isOverriden}}(m)\}$;
          $\triangleright$ Iterate over all callsites.
    22       foreach $\textbf{m}\in S$ do
    23             foreach $cs\in\textnormal{{callSites}}(\textbf{m})$ do
    24                   if $\textnormal{{isConstructor}}(cs)\wedge cs.\text{type}\in\mathcal{T}_{apis}$ then
    25                         $\mathcal{M}_{c}\leftarrow\mathcal{M}_{c}\cup{cs}$;
    26
    27                  if $\textnormal{{isStatic}}(cs)\wedge cs.\text{type}\in\mathcal{T}_{apis}$ then
    28                         $\mathcal{M}_{s}\leftarrow\mathcal{M}_{s}\cup{cs}$;
    29
    30                  if $\textnormal{{receiverType}}(cs)\in\mathcal{T}_{apis}$ then
    31                         $\mathcal{M}_{a}\leftarrow\mathcal{M}_{a}\cup{cs}$;
    32
    33      return $\mathcal{M}_{c},\mathcal{M}_{s},\mathcal{M}_{a}$;
    Algorithm 1 Identifying mockable fields, types, and scope
    Facilitating Mocking.  Mocking enables the creation of simulated objects that mimic the behavior of actual components. These are particularly useful for simulating external services, databases, or third-party API calls. By facilitating mocking, LLM-generated tests can attempt to verify code behavior and ensure proper component interaction without having to reflect on the full implementation details of all dependencies in the application. Our approach uses a deterministic methodology (shown in algorithm 1) to identify relevant candidates for mocking given a focal method ($\mathcal{M}_{focal}$), it’s corresponding focal class ($\mathcal{T}_{focal}$), and user-specified mockable third-party APIs ($\mathcal{T}_{api}$). The computed information is used to construct a test skeleton that is included in the test-generation prompt (shown in sections f, g, and h of fig. 3). The skeleton includes mocking-related definitions and setup code, leaving the task of filling the test inputs, sequence, and assertions to the LLM.
    • Identifying fields to be mocked. First, we discover all the candidate fields ($\mathcal{T}_{m}$) and types ($\mathcal{T}_{m}$) in the application that need to be mocked. To determine the candidate fields, we iterate through the fields defined in the focal class, identifying each field whose type occurs in $\mathcal{T}_{api}$ (lines 3–5). To identify all types that need to be mocked, we begin by examining the formal parameter types of the focal method $f_{m}$ and the focal class $f_{c}$ (line 6). We then perform a transitive search (lines 7–11), iterating through each type to determine whether it belongs to $\mathcal{T}_{api}$. For each identified type, we also consider the parameter types of its constructors. This process ensures that all relevant types are included. The identified fields and types to be mocked are used to populate section f of the test-generation prompt (fig. 3).
    • Identifying methods to be stubbed. Next, we identify the scope for creating stubs for mocking. These stubs emulate method calls with “when” and “then” clauses. The “when” clauses define the conditions under which a mock should return a specified value and the “then” clauses specify the expected behavior once the conditions are met. Our approach for determining this scope is shown in Function IdentifyMockingScope of algorithm 1: we start by adding the focal method ($f_{m}$) and its class constructors to the scope set. We then include methods reachable from $f_{m}$ within the focal class. If $f_{c}$ is a service entry class, we also add its overridden methods. We iterate over all methods in the scope set, examining their call sites to classify them as mockable constructor calls, static calls, or API calls based on their types. Information about these classes of calls are used to populate parts g and h of the test-generation prompt (fig. 3).

### v3

# ASTER: Natural and Multi-language Unit Test Generation with LLMs

@@ Abstract
Implementing automated unit tests is an important but time-consuming activity in software development. To assist developers in this task, many techniques for automating unit test generation have been developed. However, despite this effort, usable tools exist for very few programming languages. Moreover, studies have found that automatically generated tests suffer poor readability and do not resemble developer-written tests. In this work, we present a rigorous investigation of how large language models (LLMs) can help bridge the gap. We describe a generic pipeline that incorporates static analysis to guide LLMs in generating compilable and high-coverage test cases. We illustrate how the pipeline can be applied to different programming languages, specifically Java and Python, and to complex software requiring environment mocking. We conducted an empirical study to assess the quality of the generated tests in terms of code coverage and test naturalness—evaluating them on standard as well as enterprise Java applications and a large Python benchmark. Our results demonstrate that LLM-based test generation, when guided by static analysis, can be competitive with, and even outperform, state-of-the-art test-generation techniques in coverage achieved while also producing considerably more natural test cases that developers find easy to understand. We also present the results of a user study, conducted with 161 professional developers, that highlights the naturalness characteristics of the tests generated by our approach.
I Introduction
@@
Unit testing is a key activity in software development that serves as the first line of defense against the introduction of software bugs. Manually writing high-coverage unit tests can be tedious and time consuming. To address this, many automated test generation (ATG) techniques have been developed, aimed at reducing the cost of manual test suite development: over the last few decades, this research has produced a variety of approaches based on symbolic analysis (e.g., [1, 2, 3, 4, 5, 6, 7]), search-based techniques (e.g., [8, 9, 10, 11, 12, 13]), random and adaptive-random techniques (e.g., [14, 15, 16, 17, 18, 19]), etc.
Figure 1: Results of the survey question on whether developers would add automatically generated tests to regression test suites.
@@
These techniques have achieved considerable success in generating high-coverage test suites with good fault-detection ability, but they still have several key limitations—with respect to test readability, test scenarios covered, and test assertions created. Previous studies (e.g., [20]) have shown that developers find automatically generated tests lacking in these characteristics, suffering poor readability and comprehensibility, covering uninteresting sequences, and containing trivial or ineffective assertions. Automatically generated tests are also known to contain anti-patterns or test smells [21] and generally not perceived as being “natural” in the sense that they do not resemble the kinds of tests that developers write. All these factors inhibit the adoption of ATG tools in practice, as developers consider the tests generated by these tools to be hard to maintain and are reluctant to add them to regression test suites without some or considerable rewrite.
To understand developer perception of tests created by ATG tools, we conducted a survey of professional software developers (details in §LABEL:subsec:rq2naturalness). fig. 1 shows the results for one of the survey questions—on readiness of automatically generated tests (by two ATG tools, EvoSuite [22] for Java and CodaMosa [23] for Python) for addition to regression test suites, comparing them with human-written tests. For EvoSuite, 42% of the responses to this question stated that the tests were not suitable for addition to a regression test suite, with an additional 21% stating that the tests could be added only after significant modifications. For CodaMosa, the result is worse, with 87% of the responses falling in these two categories. These values are significantly lower for human-written tests, which shows that ATG tools need considerable improvement in this respect—so that they generate tests that are readily usable by developers in practice.
In this work, we investigate the potential of LLMs in generating natural test cases that developers consider to be readily usable for building regression test suites. Our goal is to leverage the inherent ability of LLMs in synthesizing natural-looking code, while also avoiding the pitfalls of using LLMs off-the-shelf for test generation, which can result in the creation of tests that often have compilation errors or achieve low code coverage. Moreover, by leveraging LLMs’ intrinsic knowledge of different programming languages (PL) and frameworks, our broader goal is to build a multi-language ATG tool that performs well on enterprise-grade applications, with their complex structure, framework dependencies, and multi-tiered architecture. Finally, we aim to develop a technique in which the LLM is a pluggable component and, therefore, can be configured to work with different language models.
@@
Figure 2: Illustration of naturalness (in terms of test names, variable names, and assertions) and mocking in test cases generated by the LLM-assisted technique of aster (right) compared with tests generated by EvoSuite [11] and CodaMosa [24] (left).
@@
We present a technique for LLM-assisted test generation guided by program analysis. The technique consists of preprocessing and postprocessing phases that wrap LLM interactions. The preprocessing phase performs static program analysis to compute relevant information to be included in LLM prompts for a given method under test (or focal method). This ensures that the LLM prompt has sufficient context (similar to the information that a developer would use while writing test cases for a method) and increases the chances that it generates compilable and meaningful test cases. The postprocessing phase checks the generated tests for compilation and runtime errors and constructs new prompts aimed at fixing the errors. After the test-repair step, the technique produces a set of passing tests and a set of failing tests for the focal method. To increase code coverage, the technique includes a coverage-augmentation phase, in which prompts are crafted for instructing the LLM to generate test cases aimed at exercising uncovered lines of code.
@@
We implemented the technique in a tool, called aster, for two programming languages, Java and Python, thus demonstrating the feasibility of building multilingual unit test generators with LLMs guided by lightweight program analysis. aster also incorporates mocking capability for Java unit test generation, which makes it applicable to applications that perform database operations, implement services, or use complex libraries. We present a generic approach for generation of test with mocks that is extensible to different library APIs.
@@
Results-at-a-glance. We performed a comprehensive evaluation of aster, using six general-purpose and code LLMs, to assess the generated tests in terms of code coverage and test naturalness; we also compared aster against state-of-the-art Java and Python unit test generators, EvoSuite [22] and CodaMosa [23]. Our results show that, in terms of code coverage, aster is very competitive with EvoSuite (+2.0%, -0.5%, and +5.1% line, branch, and method coverage) for Java SE applications while performing significantly better (+26.4%, +10.6%, and +18.5%, line, branch, and method coverage) for Java EE applications. For Python, aster outperforms CodaMosa considerably (+9.8%, +26.5%, and +22.5% line, branch, and method coverage). With respect to test naturalness, our quantitative evaluation shows that aster-generated tests have superior naturalness characteristics than EvoSuite- and CodaMosa-generated tests, while our user survey (of 161 professional software engineers) confirms strong developer preference for aster test cases over the tests generated by the other tools: e.g., 70% and 88% of the responses for Java and Python, respectively, stated that aster-generated tests could be added to a regression suite with no or only minor changes.
Our work corroborates and strengthens the findings on high developer acceptance rates of LLM-generated tests reported in recent industry studies  [25] by providing additional evidence in a different industry setting. Our work also provides new insights via a more rigorous evaluation of LLM-based test generation, showing that smaller models, when provided with suitable code context, can be made competitive to larger models and even outperform those models.
aster’s Java test generation capability is offered in the IBM watsonx Code Assistant for Enterprise Java Applications product [26]. The results of our experiments and the naturalness evaluator are available in our artifact [27].
@@
  • An LLM-assisted test-generation technique, and its implementation in a tool called aster, that is fueled by static analysis and can be extended to multiple PLs.
  • Empirical results showing that aster for Java and Python (configured with six different LLMs) is competitive with, and even outperforms, state-of-the-art ATG tools in code coverage achieved while generating considerably more natural tests than those tools.
@@
  • In-depth quantitative and qualitative analysis of test naturalness via an automated approach and a survey of 161 professional developers highlighting the naturalness attributes of aster-generated tests.
II Motivation
@@
The primary motivation for LLM-assisted test generation is to overcome the limitation of conventional ATG tools w.r.t lack of naturalness in the tests they create. By leveraging the LLM’s inherent ability of creating natural-looking code, we can generate more readable, comprehensible, and meaningful test cases. Another motivation is to build multi-lingual unit test generators—leveraging the LLM’s understanding of the syntax and semantics of multiple PLs on which the models are typically trained. Building such test generators using conventional approaches (e.g., symbolic or evolutionary techniques) can be challenging and no such test generators exist. In contrast, with lightweight static analysis guiding LLM interactions, a multi-lingual unit test generator can be easily implemented. Finally, an LLM-based approach can also address test generation for complex applications, which often require mocking.
@@
To illustrate these benefits, fig. 2 presents sample aster-generated test cases and tests generated by two existing ATG tools: EvoSuite [11] for Java and CodaMosa [24] for Python. The aster-generated test cases have more meaningful test names  1, variable names  2, and assertions  3 than the EvoSuite- or CodaMosa-generated tests. For instance, consider the test cases for Apache Commons CLI [28]. The variable storing the return value from flatten() is called flattenedArguments in the aster test—clearly capturing the meaning of the stored data—whereas the corresponding variable in the EvoSuite test is called stringArray1 (line 5), which captures simply the data structure type instead of any meaning of the stored data. Similarly, the Python test case generated by aster, shown in fig. 2(c), has meaningful test name and variable names. Moreover, the assertion in the test case is generated taking into account the expected transformation of the input by the focal method, converting the input hour value to minutes. fig. 2(b) shows a unit test case with mocking of library APIs generated for Apache Commons JXPath [29]. The test mocks the behavior of the org.w3c.dom.Node library class  4.
III Our Approach
Figure 3 illustrates our LLM-based test-generation technique, guided by program analysis. The process consists of two phases, both utilizing stage-specific prompting: (1) Preprocessing, in which static analysis extracts context and constructs initial prompts (fig. 4) for LLM-driven unit test generation in Java and Python and (2) Postprocessing, where the generated tests are repaired for compilation and runtime errors, and augmented for coverage enhancement.
@@ III Our Approach
Preprocessing
@@
The preprocessing phase performs static analysis of the application to collect relevant context that an LLM might require to generate unit tests. The key objective of this stage is to gather all the necessary information pertaining to the focal method and its broader context within the application. This information is used to populate fields 1 and 2 of the prompt template shown in Figure 4. In this section, we discuss the preprocessing steps for Java and how those steps help construct a prompt for the LLM.
@@
Figure 4: Templates for composing prompts for test generation, test repair, and coverage augmentation.
@@
    Testing scope. The first step in the preprocessing phase identifies the testing scope given a Java focal class $f_{c}$. The testing scope lists the set of focal methods to be targeted for test generation. This set consists of (1) public, protected, and package-visibility methods declared in $f_{c}$ and (2) any inherited method implementations from an abstract super class of $f_{c}$. The set excludes inherited methods from a non-abstract super class as those methods are targeted for test generation in the context of their declaring class as the focal class. If $f_{c}$ is an abstract class, the testing scope consists of static methods, if any, declared in $f_{c}$. The focal class and method are used to populate part d of the test-generation prompt.
@@
    Relevant constructors. aster next identifies the relevant constructors for a focal method so that their signatures can be specified to the LLM, enabling it to create required objects for invoking the focal method. The relevant constructors for a focal method $m$ include the constructors of the focal class (if $m$ is a virtual method), along with the constructors of the each formal parameter type of $m$, considering application types only (i.e., ignoring library types). The analysis is done transitively for the formal parameter types of each identified constructor, thus ensuring that the LLM prompt includes comprehensive context information about how to create instances of application types that may be needed for testing $m$. The discovered constructors are used in the test-generation prompt (fig. 4  a).
@@
    Relevant auxiliary methods. Accessor methods, which consist of getters and setters, provide a mechanism for reading and modifying internal object state, while preserving encapsulation. Our approach identifies setter methods in the focal class and in each formal parameter type of the focal method (if the type is an application class). Signatures of these setters are added to the LLM prompt to help with setting up suitable object states for invoking the focal method. For reading object state, our approach computes getters of the focal class and the return type of the focal method (limiting to application return types), and includes their signatures in the LLM prompt. This can help the LLM generate suitable assertion statements for verifying relevant object state (of the receiver object or the returned object) after the focal method call. This information is used to fill section b of the test-generation prompt in fig. 4.
@@
    Private methods.  Private methods are inaccessible outside the class. One way to test them is by using reflection (in the case of Java) to overwrite method accessibility. However, that is considered a bad practice. A more accepted approach is to test private methods indirectly by invoking them through non-private methods of the class. To facilitate this, we compute the class call graph of the focal class, identify call chains from non-private methods to private methods, and provide these call chains to the LLM to enable it to generate test cases that invoke private methods through externally visible methods. This information is used to populate  fig. 4 e section.
    Facilitating Mocking.  Mocking is a technique in software testing that allows developers to create simulated objects mimicking the behavior of actual components. This is particularly useful for simulating external services, databases, or third-party API calls. By incorporating mocking capabilities, LLM-generated tests can verify code behavior and ensure proper component interaction without needing to understand the full implementation details of all dependencies in the application. Our approach uses a systematic method to identify relevant candidates for mocking. The mocking facilitation process consists of two main steps: identifying fields and types to be mocked, and identifying methods to be stubbed. In the first step, we examine the focal class and method to discover all candidate fields and types that need to be mocked. We begin by iterating through the fields defined in the focal class, identifying each field whose type matches one of the mockable APIs. We then perform a comprehensive search starting from the formal parameter types of the focal method and class. This search is transitive, meaning we also consider the parameter types of constructors for each identified type. The identified fields and types to be mocked are used to populate section f of the test-generation prompt (fig. 4). The second step involves determining the scope for creating method stubs. These stubs emulate method calls using “when” and “then” clauses. “When” clauses define the conditions under which a mock should return a specified value, while “then” clauses specify the expected behavior once these conditions are met. To identify the stubbing scope, we start with the focal method and its class constructors. We then include all methods reachable from the focal method within the focal class. If the focal class is a service entry class, we also include its overridden methods. Finally, we examine all call sites within this scope, categorizing them as mockable constructor calls, static calls, or API calls based on their types. These information are used to populate parts g and h of the test-generation prompt (fig. 4).